# Extended-Use Designs on Very Large Online Platforms

### v1

@@ Abstract.
In the attention economy, online platforms are incentivized to maximize user engagement through extended-use designs (EUDs), even when such practices conflict with users’ best interests. We conducted a structured content analysis of all Very Large Online Platforms (VLOPs) to identify the EUDs these influential apps and sites use. We conducted this analysis posing as a teenager to understand the EUDs that young people are exposed to. We find that VLOPs use four strategies to promote extended use: pressuring, enticing, trapping, and lulling users. We report on a hierarchical taxonomy organizing the 63 designs that fall under these categories. Applying this taxonomy to all 17 VLOPs, we identify 583 instances of EUDs, with social media platforms using twice as many EUDs as other VLOPs. We present three vignettes illustrating how these designs reinforce one another in practice. We further contribute a graphical dataset of videos illustrating these features in the wild.
digital wellbeing, technology overuse, attention
@@ 1. Introduction
In the attention economy (Davenport and Beck, 2001), online platforms compete for their users’ time (Kane, 2019). They are incentivized to try to: 1) draw people in as often as possible, 2) keep them engaged for as long as possible, and 3) disclose as much as possible about their habits and preferences while they are there (see Figure 1). This combined approach maximizes the platform’s ability to serve advertisements and to make these advertisements as personalized and targeted as possible  (Lambrecht and Tucker, 2013). To guide a product toward this desired end-state, companies use engagement metrics like time-on-task, click-through rates, platform traffic, and session duration as primary indicators of their product’s success. They employ extensive A/B testing to engineer interfaces that lead to increases in these metrics (Richards and Hartzog, 2024). We refer to the resulting interface decisions as extended-use designs (EUDs).
@@
EUDs do not always align with users’ best interests. Hartford and Stein explain that digital platforms’ aggressive—and often effective—efforts to co-opt users’ attention reduce people’s autonomy, displace alternative activities they value, and incrementally reshape their life experiences (Hartford et al., 2022). Richards and Hartzog argue that the practice of optimizing for user engagement is a form of exploitation that has long flown under the radar and deserves greater scrutiny, calling it a “distinct and dangerous concept that should be specifically regulated” (Richards and Hartzog, 2024).
@@
This claim is reflective of an increasing societal and academic interest in understanding and curtailing EUDs (Mildner et al., 2023; Monge Roffarello et al., 2023; Monge Roffarello and De Russis, 2022; Docherty and Biega, 2022; Lukoff et al., 2021). And these concerns are heightened with respect to platforms’ effects on young people (of Health and Services, 2023; Woods and Scott, 2016; Viner et al., 2019; Mcculloh and Cohen, 2023). In both North America and Europe, law suits (e.g., (General, 2023)), proposed legislation (e.g., (of the Governor of New York, 2023)), and active legislation (e.g., (Senate, 2024)) all express frustration with companies’ attempts to monopolize teenagers’ time and attention. The European Parliament’s Internal Market and Consumer Protection Committee recently voted in favor of developing new regulations to prohibit EUDs, with particular attention to designs that affect minors. And in the United States, the U.S White House Interagency Kids Online Health and Safety Taskforce released design recommendations that call for reducing and removing features that encourage excessive or problematic use by youth  (Substance Abuse and Mental Health Services Administration (2021), SAMHSA).
@@
A growing body of academic literature seeks to characterize EUDs and their consequences. For example, Monge Roffarello and colleagues report on a meta-analysis of existing literature, surfacing 11 different UI patterns that attempt to capture users’ attention (Monge Roffarello et al., 2023). Flayelle and colleagues demonstrate that UI designs can systematically both: 1) exploit reflexive, automatic responses and 2) shape users’ conscious decision-making, to encourage habitual usage patterns (Flayelle et al., 2023). Gray and colleagues wrangle 10 taxonomies of 262 manipulative design patterns—including many EUDs—into a hierarchical ecology described by a closed grammar (Gray et al., 2024).
@@
In this paper, we contribute to these ongoing efforts in two ways. First, we seek to characterize the extent to which users in general and young people in particular encounter EUDs. The European Commission  [^1] The governmental body in the European Union that proposes and enforces legislation has defined “Very Large Online Platforms,” or “VLOPs,” as those which have at least 45 million monthly active users in the European Union (Commission, 2024). The designs adopted by VLOPs are poised to be the ones to which users are most exposed, given the broad reach of these platforms. The platforms with the widest adoption in the EU (Commission, 2023) are largely the ones with the widest adoption globally (Statista, 2024) and also the platforms which have the largest reach with teens (Anderson et al., 2023). Thus, we conducted a systematic evaluation of EUDs in VLOP interfaces as a way to identify the designs with the widest global impact.
@@
Secondly, we sought to examine these common patterns for themes in the strategies behind them. Prior work describes and clusters many discrete design patterns which have—and will likely continue to—change with time. As of yet, these patterns are not organized by the deeper strategies which motivate them and are more likely to persist with time, providing a more stable target for regulatory action. Specifically we asked:
  • RQ1: What EUDs are present on VLOPs?
@@
To investigate these questions, we conducted a two-part evaluation of all $N=16$ VLOP apps and $N=17$ VLOP websites to identify, characterize, and quantify the EUDs present on these platforms. We created accounts as teen users located in Europe to ensure the designs we encountered were ones that would be shown to teens. We conducted a thematic analysis of these interfaces to create a catalog of the EUDs they employed and then conducted a structured content analysis to more precisely understand their prevalence across this set of influential platforms.
We find that VLOPs use four broad strategies to promote extended use. These include designs that: entice users into engaging (i.e., using pleasure-, curiosity-, or reward-driven mechanisms to encourage a usage session), pressure users into engaging (i.e., creating social or other obligations that encourage a usage session), lull users into mindlessly continuing their interactions (i.e., using low-friction and autonomy-reducing designs that extend their usage session), and trap users by erecting barriers to disengaging (i.e., using overwhelming or obfuscating designs that reduce opportunities to discontinue their usage sessions). Platforms were most likely to use a combination of strategies, with creating chain reactions, bombarding users via multiple lines of attack, and fusing EUDs being the most common.
In this paper, we showcase exemplars of each category, describe the hierarchy of patterns we encountered, report on the prevalence of each pattern, and describe three usage vignettes that illustrate how these patterns work together to promote extended use. We further contribute a video dataset of 583 interactive instances of the 14 patterns used by VLOPs. We hope that the examples, prevalence data, and organizational scheme presented here will enable designers, researchers, and regulators to determine whether and how platforms should encourage engagement.
@@ 2. Background
2.1. The Attention Economy and Extended-Use Design
People’s struggles with online platform are a direct result of manipulative designs, incentivized by the attention economy, which attempt to extend users’ engagement with the platform as much as possible (Davenport and Beck, 2001). Manipulative designs that benefit a platform at the user’s expense (sometimes referred to as “dark patterns”  (Mathur et al., 2021)) are not unique to attention-capture; they are also used extensively to attempt to influence users’ purchase decisions  (Moser et al., 2019) and data disclosure  (Zhang-Kennedy et al., 2024). Historically, research on manipulative designs has focused on financial harms, studying eCommerce platforms that exert purchase pressure by, for example, enrolling users in hidden subscriptions or adding items to their shopping cart without consent (e.g.,  (Bongard-Blanchy et al., 2021; Moser et al., 2019; Mathur et al., 2019; Luguri and Strahilevitz, 2021)). However, more recent work has begun to characterize dark patterns that manipulate users into extended engagement. There is a pressing need to understand these manipulative EUDs, because they undermine users’ attempts to cultivate a healthy relationship with technology and set boundaries that allow them to direct their attention as needed toward sleep, work, school, time with loved ones, time outdoors, and other activities that are conducive to emotional and physical wellbeing.
@@ 2.1. The Attention Economy and Extended-Use Design
There is also increasing public interest in understanding and regulating manipulative EUDs. Manipulative designs more broadly, including those that lead to financial harms or compromise data privacy, are already the target of regulatory action, and for example, the Digital Services Act Article 25 states that “providers of online platforms shall not design, organise or operate their online interfaces in a way that deceives or manipulates the recipients of their service or in a way that otherwise materially distorts or impairs the ability of the recipients of their service to make free and informed decisions.” In a recent report to the European Parliament’s Committee on Internal Market and Consumer Protection, Van Sparrentak called for further legislation to address “addictive” EUDs in particular. This call to action lists a wide variety of manipulative practices, including quantification of social approval, overload of sensory stimuli, limited control over interfaces, notifications demanding a user return to an app on a fixed schedule, video autoplay, and pull-to-refresh content displays (Monge Roffarello et al., 2023).
@@
However, no prior work has documented the prevalence of EUDs on technology platforms. Regulators also need a reproducible taxonomy for identifying common EUDs, as well as the underlying psychological mechanisms being manipulated, to move forward with policy making and enforcement. To this end, we reviewed existing taxonomies of manipulative “dark pattern” designs and theories of online behavioral addictions, and then leveraged these theories and empirical review of VLOPs to create a novel and reproducible taxonomy.
2.2. Theorizing Extended-Use Design
Organizational taxonomies describing categories of manipulative designs have been described either through empirical review of digital products or theoretically driven frameworks. For example, Zagal et al.’s description of dark patterns in video game design categorized them into: 1) temporal dark patterns (grinding, playing by appointment), 2) monetary dark patterns (pay to skip, pre-delivered content, monetized rivalries), and 3) social capital-based dark patterns (social pyramid schemes, impersonation)  (Zagal et al., 2013). Taxonomies of dark patterns in e-commerce platforms have focused on purchase pressure and deceptive practices  (Moser et al., 2019), while an analysis of children’s mobile games described additional manipulative features such as parasocial relationship pressure and navigational constraints (Radesky et al., 2022).
Most relevant to the interrogation of EUDs are taxonomies of attention-capture designs, which encourage users to initiate or extend engagement with a platform. Monge Roffarello and colleagues conducted a systematic literature review of 43 scholarly articles that define dark patterns and then used this metaanalysis to create a typology of 11 common attention-capture patterns (see Table 1) that: 1) exploit known psychological biases and heuristics, 2) remove the need for decision-making, 3) lead the user to lose track of goals, 4) lead to a lost sense of time and control, and 5) lead to a sense of regret about the time spent on a digital service.
@@ 2.2. Theorizing Extended-Use Design
In a complementary approach, Flayelle and colleagues (2023) used a theoretical framework to classify the categories of designs that could contribute to addictive online behaviors such as video gaming, online gambling, cybersexual activities, online shopping, social networking, and binge-watching  (Flayelle et al., 2023). Their taxonomy split design features into “model-free” and “model-based” mechanisms. Model-free mechanisms do not depend on users developing a mental model to motivate behavior and thus only depend upon operant conditioning to build habits through experience, strengthen those habits through reinforcement learning, and develop automatic tendencies that cause the user to reflexively engage with a platform in response to cues or triggers. Such mechanisms are likely to be more powerful in children than in adults, as young children are more sensitive to behavioral reinforcement, are attracted to novel cues, and have not yet developed the metacognitive abilities that enable abstract decision-making. In contrast, model-based mechanisms enable users to exert control by comparing the benefits and costs of potentially problematic activity with alternatives based on the subjective value of each. Digital platforms can exploit model-based mechanisms by influencing how the user over-values an online experience, reducing the user’s deliberation or interfering with decision-making, leveraging cognitive biases, and allowing only partial goal fulfillment so that the user feels compelled to keep engaging. Again, children and teens are more likely to be influenced by model-based mechanisms because their decision-making is more likely to be influenced by automatic reflexes, with less ability to delay gratification when faced with pleasurable online experiences, and lower planning/organization skills that allow users to compare the costs and benefits of a decision.
While existing taxonomies have explored design categories that could extend user engagement, they have focused more narrowly on individual domains like social media or gaming. To fill this gap, we examined the designs present on all VLOPs in hopes of identifying the extended use design machenisms that have the widest global impact, and moving beyond the scope of prior examinations.
2.3. Adolescent Development and Manipulative Design

### v2

# The Engagement-Prolonging Designs Teens Encounter on Very Large Online Platforms

@@ Abstract.
In the attention economy, online platforms are incentivized to design products that maximize user engagement, even when such practices conflict with users’ best interests. We conducted a structured content analysis of all Very Large Online Platforms (VLOPs) to identify the designs these influential apps and sites use to capture attention and extend engagement. Specifically, we conducted this analysis posing as a teenager to identify the designs that young people are exposed to. We find that VLOPs use four strategies to extend teens’ use: pressuring, enticing, trapping, and lulling them into spending more time online. We report on a hierarchical taxonomy organizing the 63 designs that fall under these categories. Applying this taxonomy to all 17 VLOPs, we identify 583 instances of engagement-prolonging designs, with social media platforms using twice as many as other VLOPs. We present three vignettes illustrating how these designs reinforce one another in practice. We further contribute a graphical dataset of videos illustrating these features in the wild.
digital wellbeing, ethical interfaces, teens, technology overuse, manipulative design
@@ 1. Introduction
In the attention economy (Davenport and Beck, 2001), online platforms compete for their users’ time and attention (Kane, 2019). They are incentivized to try to: 1) draw people in as often as possible, 2) keep them engaged for as long as possible, and 3) encourage them to disclose as much as possible about their habits and preferences while they are there (see Figure 1). This combined approach maximizes the platform’s ability to serve advertisements and to make these advertisements as personalized as possible  (Lambrecht and Tucker, 2013). To guide a product toward this desired end-state, companies use engagement metrics like time-on-task, click-through rates, platform traffic, and session duration as primary indicators of their product’s success. They employ extensive A/B testing to engineer interfaces that lead to increases in these metrics (Richards and Hartzog, 2024). We refer to the designs that extend the time people spend on the platform, heighten the frequency with which they interact with the platform, and deepen their disclosures to the platform and third parties, as engagement-prolonging designs (EPDs).
@@
EPDs do not always align with users’ best interests. Hartford and Stein explain that digital platforms’ aggressive—and often effective—efforts to co-opt users’ attention reduce people’s autonomy, displace alternative activities they value, and incrementally reshape their life experiences (Hartford et al., 2022). Richards and Hartzog argue that the practice of optimizing for user engagement is a form of exploitation that has long flown under the radar and deserves greater scrutiny, calling it a “distinct and dangerous concept that should be specifically regulated” (Richards and Hartzog, 2024).
This claim is reflective of an increasing societal and academic interest in understanding and curtailing the practice of designing to prolong engagement (Mildner et al., 2023; Monge Roffarello et al., 2023; Monge Roffarello and De Russis, 2022; Docherty and Biega, 2022; Lukoff et al., 2021). And these concerns are heightened with respect to platforms’ effects on youth (of Health and Services, 2023; Woods and Scott, 2016; Viner et al., 2019; Mcculloh and Cohen, 2023). Relative to adults, children and teens have less inhibitory control (Casey et al., 2008; Posner and Rothbart, 2000), making them more vulnerable to manipulative designs that erode self-control. They also more likely to adopt new technologies, relative to the adults in their lives (Lee, 2007), and their social worlds are more likely to have important online dimensions (Boyd, 2014), making it essential to create respectful online spaces where young people are free from manipulation by powerful companies seeking to monetize their attention.
@@
In both North America and Europe, law suits (e.g., (General, 2023)), proposed legislation (e.g., (of the Governor of New York, 2023)), and active legislation (e.g., (Senate, 2024)) all reflect frustration with companies’ attempts to prolong teenagers’ engagement with their products. The European Parliament’s Internal Market and Consumer Protection Committee recently voted in favor of developing new regulations to prohibit EPDs, with particular attention to designs that affect minors. And in the United States, the U.S White House Interagency Kids Online Health and Safety Taskforce released design recommendations that call for reducing and removing features that encourage excessive or problematic use by youth (Substance Abuse and Mental Health Services Administration (2021), SAMHSA).
@@
A growing body of academic literature seeks to characterize EPDs and their consequences. For example, Monge Roffarello and colleagues report on a meta-analysis of existing literature, surfacing 11 different UI patterns that use deceptive techniques to capture users’ attention (Monge Roffarello et al., 2023). Flayelle and colleagues demonstrate that UI designs can systematically both: 1) exploit reflexive, automatic responses and 2) shape users’ conscious decision-making, to encourage habitual usage patterns (Flayelle et al., 2023). Gray and colleagues wrangle 10 taxonomies of 262 manipulative design patterns—including many EPDs—into a hierarchical ecology described by a closed grammar (Gray et al., 2024). These prior efforts both reflect the HCI community’s interest in characterizing this design space and demonstrate the complexity, diversity, and ubiquity of EPDs.
@@
In this paper, we contribute to these ongoing efforts in two ways. First, we seek to characterize the extent to which young people in particular encounter EPDs. The European Commission  [^1] The governmental body in the European Union that proposes and enforces legislation has defined “Very Large Online Platforms,” or “VLOPs,” as those which have at least 45 million monthly active users in the European Union (Commission, 2024). The designs adopted by VLOPs are poised to be the ones to which users are most exposed, given the broad reach of these platforms and the fact that many smaller companies copy patterns established by VLOPs. The platforms with the widest adoption in the EU (Commission, 2023) are largely the ones with the widest adoption globally (Statista, 2024) and also the platforms which have the largest reach with teens (Anderson et al., 2023). Thus, we conducted a systematic evaluation of EPDs in VLOP interfaces as a way to identify the designs with the widest global impact on teen users.
@@
Second, we sought to examine these common patterns for themes in the strategies behind them. Prior work describes and clusters many discrete design patterns which have—and will likely continue to—change with time. As of yet, these patterns are not organized by the deeper strategies which motivate them and are more likely to persist even as design techniques change, providing a more stable target for regulatory action. Specifically we asked:
  • RQ1: What EPDs do VLOPs present to teen users?
@@
To investigate these questions, we conducted a two-part evaluation of all $N=16$ VLOP apps and $N=17$ VLOP websites to identify, characterize, and quantify the EPDs present on these platforms. We created accounts as teen users located in Europe to ensure the designs we encountered were ones that would be shown to teens. We conducted a thematic analysis of these interfaces to create a catalog of the EPDs they employed and then conducted a structured content analysis to more precisely understand their prevalence across this set of influential platforms.
We find that VLOPs use four broad strategies to promote extended use. These include designs that: entice users into engaging (i.e., using pleasure-, curiosity-, or reward-driven mechanisms to encourage a usage session), pressure users into engaging (i.e., creating social or other obligations that encourage a usage session), lull users into mindlessly continuing their interactions (i.e., using low-friction and autonomy-reducing designs that extend their usage session), and trap users by erecting barriers to disengaging (i.e., using overwhelming or obfuscating designs that reduce opportunities to discontinue their usage sessions). Platforms were most likely to use a combination of strategies.
In this paper, we showcase exemplars of each category, describe the hierarchy of patterns we encountered, report on the prevalence of each pattern, and describe three usage vignettes that illustrate how these patterns work together to prolong engagement. We further contribute a video dataset of 583 interactive instances of the 14 patterns used by VLOPs. We hope that the examples, prevalence data, and organizational scheme presented here will enable designers, researchers, and regulators to determine whether and how platforms should encourage teen users to engage with a platform.
@@ 2. Background
2.1. Manipulative Designs
Deceptive designs that benefit a platform at the user’s expense have proliferated across digital products (Mathur et al., 2021). These designs have historically been referred to as “dark patterns,” a term that has been increasingly been replaced with “deceptive designs” (Brignull, 2020) or “manipulative designs” (Radesky et al., 2022) to move away from racialized language that uses “dark” to convey a sense of danger or maliciousness. User Experience practitioner, Harry Brignull, was the first to call widespread attention to the practice of manipulative design. Using examples from shopping and travel websites, Brignull created a taxonomy documenting patterns that deceive users into taking actions against their own intentions or best interest (such as the “Bait and Switch” patterns, in which the user intends to perform one action but is tricked by the interface into performing another, or the “Confirmshaming” pattern, in which the platform erects a dialog box with shaming language as a barrier to proceeding). Building on Brignull’s original collection of examples, Conti and Sobiesk created a taxonomy of malicious interface design techniques, which they defined as interfaces that manipulate, exploit, or attack users (Conti and Sobiesk, 2010). Gray et al. later presented a broader categorization of Brignull’s taxonomy and collapsed many patterns into broader categories, such as “Nagging” (repeatedly making the same request to the user) and “Obstruction” (preventing the user from accessing functionality) (Gray et al., 2018).
Many studies of manipulative designs have focused on financial harms, studying eCommerce platforms that exert purchase pressure by, for example, enrolling users in hidden subscriptions or adding items to their shopping cart without consent (e.g.,  (Bongard-Blanchy et al., 2021; Moser et al., 2019; Mathur et al., 2019; Luguri and Strahilevitz, 2021)). Other work has shown that harms are not always monetary, and Zagal et al.’s description of manipulative designs in video games categorized these patterns into: 1) temporal manipulative designs (grinding, playing by appointment), 2) monetary manipulative designs (pay to skip, pre-delivered content, monetized rivalries), and 3) social capital-based manipulative designs (social pyramid schemes, impersonation)  (Zagal et al., 2013). Many other studies have shown that manipulative designs frequently compromise users’ privacy or attempt to harvest data deceptively (Zhang-Kennedy et al., 2024; Susser et al., 2019; Bongard-Blanchy et al., 2021).
@@ 2.1. Manipulative Designs
Manipulative designs broadly, including both those that lead to financial harms and those that compromise data privacy, are already the target of regulatory action. For example, the Digital Services Act Article 25 states that “providers of online platforms shall not design, organize or operate their online interfaces in a way that deceives or manipulates the recipients of their service or in a way that otherwise materially distorts or impairs the ability of the recipients of their service to make free and informed decisions.”
2.2. Manipulative Designs and Engagement
More recently, the conversation about manipulative designs has expanded to include platforms’ attempts to prolong users’ engagement. Designs that prolong engagement might seem innocuous, or even potentially beneficial and reflective of a user enjoying their experience with a product (Wojtowicz et al., 2024). However, too often, EPDs are only effective because of the extreme power imbalance between companies that seek to maximize engagement and the people who use these products (Hartzog and Richards, 2021). Recent legal scholarship explains that engagement tactics are frequently unfair, deceptive, or abusive, and disproportionately benefit the party stimulating the engagement while burdening the engaged (Richards and Hartzog, 2024). There is a pressing need to understand these manipulative EPDs, because they undermine users’ attempts to cultivate a healthy relationship with technology and set boundaries that allow them to direct their attention as needed toward sleep, work, school, time with loved ones, time outdoors, and other activities that are conducive to emotional and physical wellbeing.
Thus, recent scholarship in human-computer interaction has begun to characterize manipulative designs that target users’ attention (in contrast to earlier studies which focused on manipulative designs targeting financial deception). Monge Roffarello and colleagues conducted a systematic literature review of 43 scholarly articles (Monge Roffarello et al., 2023) that identify attention-related manipulative designs and then used this meta-analysis to create a typology of 11 common attention-capture patterns (see Table 1) that: 1) exploit known psychological biases and heuristics, 2) remove the need for decision-making, 3) lead the user to lose track of goals, 4) lead to a lost sense of time and control, and 5) lead to a sense of regret about the time spent on a digital service. They divided the 11 attention-capture patterns into two categories: deceptive designs that trick the user into a false belief, such as Time Fog and Fake Social Notifications, and seductive designs that tempt the user with short-term satisfaction like Guilty Pleasure Recommendations. Similarly, Flayelle and colleagues used a theoretical framework to classify the categories of designs that could contribute to addictive online behaviors such as video gaming, online gambling, cybersexual activities, online shopping, social networking, and binge-watching  (Flayelle et al., 2023).
There is increasing public interest in understanding and regulating manipulative EPDs. In a recent report to the European Parliament’s Committee on Internal Market and Consumer Protection, Van Sparrentak called for further legislation to address “addictive” EPDs in particular. This call to action lists a wide variety of manipulative practices, including quantification of social approval, overload of sensory stimuli, limited control over interfaces, notifications demanding a user return to an app on a fixed schedule, video autoplay, and pull-to-refresh content displays (Monge Roffarello et al., 2023).
@@ 2.2. Manipulative Designs and Engagement
We build on this foundational prior work in several ways. Our study expands beyond attention capture and more holistically considers the ways in which platforms both trigger and prolong usage. We further characterize the prevalence of EPDs on the most widely used technology platforms (including those most frequently used by teens), and we document the EPDs that these platforms employ specifically when these interfaces are used by young people. Regulators also need a reproducible taxonomy for identifying common EPDs, as well as the underlying psychological mechanisms being manipulated, to move forward with policy making and enforcement. To this end, we reviewed existing taxonomies of manipulative designs and theories of online behavioral addictions, and then leveraged these theories and empirical review of VLOPs to create a novel and reproducible taxonomy.
@@
2.3. Adolescent Development and Engagement-Prolonging Design
A growing body of work examines the manipulative designs that child and teen users encounter. For example, Stockman and Nottingham document the use of “cuteness” as a UI tactic to manipulate both child and adult users (Stockman and Nottingham, 2024). Other work has documented the manipulative designs in games (Sousa and Oliveira, 2023) and apps (Radesky et al., 2022) used by young children, finding that these patterns are widespread in children’s products. Fitton and Read (Fitton and Read, 2019) worked with early adolescent girls to create a framework for identifying manipulative designs in free-to-play apps. An exploration of teens’ mental models of manipulative designs revealed that they are aware of widespread deceptive practices but often lack a precise understanding of when and how they are being manipulated in the moment (Renaud et al., 2024).